Use the following information to prepare the September cash budget for PTO Manufacturing Co. The following information relates to expected cash receipts and cash payments for the month ended September 30.

  1. Beginning cash balance, September 1, $44,000.
  2. Budgeted cash receipts from sales in September, $262,000.
  3. Raw materials are purchased on account. Purchase amounts are: August (actual), $72,000, and September (budgeted), $100,000. Payments for direct materials are made as follows: 65% in the month of purchase and 35% in the month following purchase.
  4. Budgeted cash payments for direct labor in September, $39,000.
  5. Budgeted depreciation expense for September, $3,600.
  6. Other cash expenses budgeted for September, $56,000.
  7. Accrued income taxes payable in September, $10,600.
  8. Bank loan interest payable in September, $1,900.
Answer #1
PTO Manufacturing co cash budget for the moth sept 30
Particulars Amount Amount
beginning cash balance sept 1 $     44,000.00
add: Cash receipts $ 2,62,000.00
Total Receipts $ 3,06,000.00
less: raw material $ 90,200.00
Direct labour $ 39,000.00
other expenses $ 56,000.00
Tax paid $ 10,600.00
Bank interest $   1,900.00
Total cash disbursement $ 1,97,700.00
Ending cash balance sept 30 $ 1,08,300.00
Working
raw material
August = 72000*35% = 25200
September = 100000*65% = 65000
Total cash paid = 25200+65000 = 90200

Give a Thumbs up if it is helpful

please let me know if any doubt